Gold-198 has a half-life of 2.7 days. How much of a 941.5 mg gold-198 sample will remain after 13.5 days?
Answer #1

K = 0.693/t1/2

t1/2 = 2.7days

K   = 0.693/2.7    = 0.256days^-1

K   = 2.303/t log[A0]/[A]

0.256 = 2.303/13.5 log941.5/[A]

0.256    = 0.17log941.5/[A]

log941.5/[A]   = 0.256/0.17

log941.5/[A]   = 1.5

   941.5/[A]   = 10^1.5

941.5/[A]      = 31.62

[A]                = 941.5/31.62    = 29.77mg

The amount of gold-198 remains after = 29.77mg >>>answer